There are some diamonds..
7 people will share diamonds equally...
If 6 members will share then 5 diamonds will be left...
If  5 members will share then 4 diamonds will be left...
If 4 members will share then 3 diamonds will be left...
If 3 members will share 2 diamonds will be left...
If 2 members will 1 diamond will be left.....
So Total how many diamonds are there
...

Grade:12th pass

1 Answers

Aditya Gupta
5 years ago
hello divya. this is a question based on common multiples. so let total diamonds be d. then d+1 will be divisible by 2, 3, 4, 5, 6 ; which also implies that d+1 will be a multiple of the LCM of these nos. moreover d should be divisible by 7.
now LCM of 2, 3, 4, 5, 6 = 60
but d+1=60 or d= 59 is not div by 7
so d+1= 2*60=120
or d= 119 which is divisible by 7.
so 119 is the smallest no of diamonds which satisfy the above conditions.
the next no of possible diamonds is 539....